OpenStudy (anonymous):

Determine whether the lines through the pairs of points are parallel, perpendicular, or neither. 3x - 2y = -5 2x + 3y = -15

OpenStudy (anonymous):

For the first equation, I can get y by itself 3x - 2y = -5 Subtract 3x from both sides -2y = -3x - 5 Divide all terms by -2 y = (3/2)x + 5/2 The slope is 3/2. Can you do the same for the second equation?

OpenStudy (anonymous):

Find the slopes of both lines by putting the equations into y=mx+b form. Then compare the two slopes. If they have the same slope - your lines are parallel. If the slopes are opposite reciprocals - your lines are perpendicular. Otherwise they're neither.